The length of a rectangle is 3 meters longer than the width. If the area is 33 square meters, find the rectangle'a dimensions.
Answer by richwmiller(17219)

len=3+wid
len*wid=33
(3+wid)*wid=33
3* wid^2=33
wid^2=11
wid=sqrt(11)
len =3+sqrt(11)
I don't like these numbers.
I suspect something is wrong with the problem